# Is it possible to describe an electron by the evolution of a single point?

### v1

@@ 2.1 The Classical Dirac Particle
like in the case of the point particle, where the factor $\gamma(v)=(1-v^{2}/c^{2})^{-1/2}$. With the definition of the CM $\bi q$}
@@
where $\bi a$} is the acceleration of the CC, the fourth order differential equations of the point $\bi r$} can be separated into a coupled system of second order differential equations for both points $\bi q$} and $\bi r$}.
@@
where the external force ${\bi F}=e{\bi E}(t,{\bi r})+e{\bi u}\times{\bi B}(t,{\bi r})$, is the Lorentz force, and $m$ and $e$ are the mass and the electric charge of the electron, respectively. The two velocities satisfy the constraints $u=c$, $v<c$. In the derivation of these dynamical equations we have made no use of a possible radiation reaction force. Equation (3) is just $d{\bi p}/dt={\bi F}$ and equation (4) is equation (2) when writing on the left hand side the CC acceleration $\bi a$}.
@@
Now the electron has two characterictic points $\bi q$} and $\bi r$} and therefore two different angular momenta can be defined with respect to both points. By the Noether analysis it is found that the angular momentum with respect to the CC takes the form:
@@
perpendicular to the velocity and acceleration of the CC, and using (2) to substitute the expression of $\bi a$}, can be rewritten as
@@
These observables represent the spin of the electron $\bi S$} and ${\bi S}_{CM}$ with respect to the CC and CM, respectively. In the center of mass frame (${\bi q}={\bi v}=0$) and they take the same value. They satisfy two different dynamical equations. In the free case $d{\bi S}/dt={\bi p}\times{\bi u}$ and $d{\bi S}_{CM}/dt=0$. The CM spin is conserved but the CC spin $\bi S$} satisfies the same dynamical equation than Dirac’s spin operator in the quantum case. It corresponds to the classical spin observable equivalent to the quantum mechanical Dirac’s spin. In the center of mass frame and for a free electron, the spin is constant and takes the value
${\bi S}=-m{\bi r}\times{\bi u},$
so that the point $\bi r$} describes a circle of constant radius $R_{0}$ at the speed of light in a plane orthogonal to the constant spin. In the center of mass frame the separation $R_{0}=|{\bi r}-{\bi q}|$ is constant and when quantizing the model takes the value $R_{0}=\hbar/2mc$, and the internal motion has an angular velocity $\omega_{0}=2mc^{2}/\hbar$, twice the frequency postulated by De Broglie. The value of the spin is $S(0)=\hbar/2$.
The angular velocity of the comoving frame $\boldsymbol{\omega}=\boldsymbol{\omega}_{p}+\boldsymbol{\omega}_{u}$, has two components, one $\boldsymbol{\omega}_{p}$ perpendicular to the plane subtended by $\bi u$} and $\bi a$} and that produces the curvature of the trajectory of the point and another $\boldsymbol{\omega}_{u}$ along the velocity $\bi u$} that gives rise to the torsion of the trajectory. They are expressed as:
@@
The last term is in fact ${\bi S}\cdot{\boldsymbol{\omega}}_{p}>0$, according to (5) and (8), because these two vectors are collinear but in opposite directions. The spin $\bi S$} is orthogonal to the velocity $\bi u$}, and therefore to $\boldsymbol{\omega}_{u}$, and this term can be rewritten as ${\bi S}\cdot{\boldsymbol{\omega}}>0$. The energy contains two terms: the first term, proportional to the linear momentum, represents the translational energy, while the second, proportional to the spin and angular velocity and to the motion of the CC (zitterbewegung), and independent of the CM velocity, represents the rotational energy $H_{rot}$ which is positive definite, because the spin $\bi S$} has the opposite direction to the angular velocity $\boldsymbol{\omega}_{p}$, and never vanishes. If we introduce in (10) the expressions of $\bi p$} in (1) and $\bi S$} in (6),
$H=\gamma(v)m{\bi v}\cdot{\bi u}+\gamma(v)m\left[({\bi r}-{\bi q})\times{\bi u}\right]\cdot\left[\frac{1}{c^{2}}{\bi u}\times{\bi a}\right],$
@@
The Classical Dirac particle has two invariant properties related to the two Casimir operators of the Poincaré grup, namely the invariant related to the four-momentum $p^{\mu}\equiv(H/c,{\bi p})$, $p^{\mu}p_{\mu}=m^{2}c^{2}$, and the invariant related to the Pauli-Lubanski four-vector $w^{\mu}=(1/2)\epsilon^{\mu\nu\sigma\lambda}p_{\nu}J_{\sigma\lambda}$, that can be written in terms of $H$, $\bi p$} and ${\bi S}_{CM}$ as:
@@
In this way the magnetic moment is $\mu_{CM}=e\hbar/2m$, Bohr’s magneton and the electric dipole $d_{CM}=e\hbar/2mc$.
@@
In this analysis no hypothesis about the size or shape of the electron has been stated and no requirements concerning a possible charge or mass distribution. The physical properties of the electron are determined by the evolution of a single point, the center of charge. It is a system of only three degrees of freedom $\bi r$}, but they satisfy a system of ordinary differential equations of fourth order and not of second order, as it corresponds to the description of the spinless point electron. In three dimensional Euclidean space the most general continuous and differentiable curve satisfies necessarily a system of ordinary fourth order diferential equations.
@@
The historical suggestion in some physics books that the fundamental equations of physics are second order differential equations represents a simplifying restriction. This requires that a Lagrangian description of a moving point will be a function of the point and its velocity, Euler-Lagrange equations are of second order and thus the trajectories of the point have no torsion, which is equivalent to the suppression of spin. It is the mechanical description of spinless objects and in the real world experimental physics has shown that there are no spinless elementary particles.

### v2

@@ 2.1 The Classical Dirac Particle
like in the case of the point particle, where the factor $\gamma(v)=(1-v^{2}/c^{2})^{-1/2}$. The definition at time $t$ of the CM $\bi q$} is
@@
where $\bi a$} is the acceleration of the CC, and all magnitudes evaluated at the same time. With this definition the fourth order differential equations of the point $\bi r$} can be separated into a coupled system of second order differential equations for both points $\bi q$} and $\bi r$}.
@@
where the external force ${\bi F}=e{\bi E}(t,{\bi r})+e{\bi u}\times{\bi B}(t,{\bi r})$, is the Lorentz force, and $m$ and $e$ are the mass and the electric charge of the electron, respectively. The two velocities satisfy the constraints $u=c$, $v<c$. In the derivation of these dynamical equations we have made no use of a possible radiation reaction force. Equation (7) is just $d{\bi p}/dt={\bi F}$, after taking the time derivative of ${\bi p}=\gamma(v)m{\bi v}$, and leaving on the left hand side the CM acceleration $d{\bi v}/dt$. The equation (8) is equation (6) when writing on the left hand side the CC acceleration $\bi a$}.
@@
Now the electron has two characteristic points $\bi q$} and $\bi r$} and therefore two different angular momenta can be defined with respect to both points. By the Noether analysis it is found that the angular momentum with respect to the CC at time $t$ takes the form:
@@
perpendicular to the velocity and acceleration of the CC. If we use the expressions of $H$ and $\bi p$} in (5) and formula (6) to substitute the expression of $\bi a$}, this spin can be rewritten as
@@
These observables represent the spin of the electron $\bi S$} and ${\bi S}_{CM}$ with respect to the CC and CM, respectively at the same time $t$. In the center of mass frame (${\bi q}={\bi v}=0$) and both spins take the same value. They satisfy two different dynamical equations. In the free case $d{\bi S}/dt={\bi p}\times{\bi u}$ and $d{\bi S}_{CM}/dt=0$. The CM spin is conserved but the CC spin $\bi S$} satisfies the same dynamical equation than Dirac’s spin operator in the quantum case. It corresponds to the classical spin observable equivalent to the quantum mechanical Dirac’s spin. In the center of mass frame and for a free electron, the spin is constant and takes the value
${\bi S}=-m{\bi r}\times{\bi u},$ (12)
so that the point $\bi r$} describes a circle of constant radius $R_{0}$ at the speed of light in a plane orthogonal to the constant spin, as depicted in figure 1.
In the center of mass frame the separation $R_{0}=|{\bi r}-{\bi q}|$ is constant and when quantizing the model takes the value $R_{0}=\hbar/2mc$, and the internal motion has an angular velocity $\omega_{0}=2mc^{2}/\hbar$, twice the frequency postulated by De Broglie. The value of the spin is $S(0)=\hbar/2$.
Figure 1: Motion of the CC of the electron $\bi r$} in the center of mass frame given in (12). The CM is located at the origin in this frame. The angular velocity $\boldsymbol{\omega}$, because the CM velocity v=0{v=0}, and according to (14) has no component along the velocity $\bi u$}, the trajectory is flat and has no torsion. The spin in the center of mass frame has the opposite direction to the angular velocity $\boldsymbol{\omega}$.
The angular velocity of the comoving frame $\boldsymbol{\omega}=\boldsymbol{\omega}_{p}+\boldsymbol{\omega}_{u}$, has two components, one $\boldsymbol{\omega}_{p}$ perpendicular to the plane subtended by $\bi u$} and $\bi a$} and that produces the curvature of the trajectory of the point and another $\boldsymbol{\omega}_{u}$ along the velocity $\bi u$}, like the Darboux vector in the Frenet-Serret formalism, that gives rise to the torsion of the trajectory. They are expressed as:
@@
In natural units, the angular velocity is related to the Darboux vector $\boldsymbol{\omega}=\tau{\bi t}+\kappa{\bi b}$, so that the torsion is the component $\omega_{u}$ along the CC velocity and the curvature $\omega_{p}$ is along the binormal. The two components of the angular velocity of the comoving rotating frame correspond to the two intrinsic geometric properties of the trajectory of the CC.
2.2 Dirac’s Hamiltonian
@@
The last term is in fact ${\bi S}\cdot{\boldsymbol{\omega}}_{p}>0$, according to (9) and (13), because these two vectors are collinear but in opposite directions, as seen in figure 1. The spin $\bi S$} is orthogonal to the velocity $\bi u$}, and therefore to $\boldsymbol{\omega}_{u}$, and this term can be rewritten as ${\bi S}\cdot{\boldsymbol{\omega}}>0$. The energy contains two terms: the first term, proportional to the linear momentum, represents the translational energy, while the second, proportional to the spin and angular velocity and to the motion of the CC (zitterbewegung), and independent of the CM velocity, represents the rotational energy $H_{rot}$ which is positive definite, because the spin $\bi S$} has the opposite direction to the angular velocity $\boldsymbol{\omega}_{p}$, and never vanishes. If we introduce in (15) the expressions of $\bi p$} in (5) and $\bi S$} in (10),
$H=\gamma(v)m{\bi v}\cdot{\bi u}-\gamma(v)m\left[({\bi r}-{\bi q})\times{\bi u}\right]\cdot\left[\frac{1}{c^{2}}{\bi a}\times{\bi u}\right],$
@@
2.3 Intrinsic properties of the Dirac particle
The Classical Dirac particle has two invariant properties related to the two Casimir operators of the Poincaré group, namely the invariant related to the four-momentum $p^{\mu}\equiv(H/c,{\bi p})$, $p^{\mu}p_{\mu}=m^{2}c^{2}$, and the invariant related to the Pauli-Lubanski four-vector $w^{\mu}=(1/2)\epsilon^{\mu\nu\sigma\lambda}p_{\nu}J_{\sigma\lambda}$, that can be written in terms of $H$, $\bi p$} and ${\bi S}_{CM}$ as:
@@ 2.3 Intrinsic properties of the Dirac particle
In this way the magnetic moment is $\mu_{CM}=e\hbar/2m$, Bohr’s magneton and the electric dipole, $d_{CM}=e\hbar/2mc$. The quantization of this particle implies that the quantum electric dipole moment operator is the operator found by Dirac in 1928, but he considered no relevant from the physical point of view, perhaps for some possible loss of spherical symmetry of the charge distribution [6]. This electric dipole gives rise to Darwin’s term of Dirac Hamiltonian.
@@
In this analysis no hypothesis about the size or shape of the electron has been stated and no requirements concerning a possible charge or mass distribution. The physical properties of the electron are determined by the evolution of a single point, the center of charge. It is a system of only three degrees of freedom $\bi r$}, but they satisfy a system of ordinary differential equations of fourth order and not of second order, as it corresponds to the description of the spinless point electron. In three dimensional Euclidean space the most general continuous and differentiable curve satisfies necessarily a system of ordinary fourth order differential equations.
@@
The historical suggestion in some physics books that the fundamental equations of physics are second order differential equations represents a simplifying restriction. This requires that a Lagrangian description of a moving point will be a function of the point and its velocity, Euler-Lagrange equations are of second order and thus the trajectories of the point have no torsion, which is equivalent to spin suppression. It represents the mechanical description of spinless particles but in the real world experimental physics has shown that there are no spinless elementary particles.
3 Natural units
The complete formalism can be done in terms of dimensionless variables once some basic properties are fixed. For the natural system of units we take as usual $\hbar=c=1$, the mass and charge of the electron $m=e=1$, the unit of length is $2R_{0}=\hbar/mc=1$, and the unit of time is $\tau_{0}=2R_{0}/c$, where $R_{0}$ is the separation between the CC and the CM for the center of mass observer. Then if we define new dimensionless position variables ${\widetilde{\bi r}}={\bi r}/2R_{0}$, ${\widetilde{\bi q}}={\bi q}/2R_{0}$, new dimensionless velocity variables ${\widetilde{\bi u}}={\bi u}/c$, ${\widetilde{\bi v}}={\bi v}/c$, and a new dimensionless time ${\widetilde{t}}={t}/\tau_{0}$, the differential equations are written in terms of dimensionless variables, and once the tildes are removed, they become:
$\frac{d{\bi q}}{dt}$ $=$ ${\bi v},\quad\frac{d{\bi v}}{dt}=\frac{1}{\gamma(v)}\left[{\bi F}-{\bi v}\left({\bi F}\cdot{\bi v}\right)\right],$ (19)
$\frac{d{\bi r}}{dt}$ $=$ ${\bi u},\quad\frac{d{\bi u}}{dt}=\frac{1-{\bi v}\cdot{\bi u}}{({\bi q}-{\bi r})^{2}}({\bi q}-{\bi r}),$ (20)
with the constraints $u=1$, $v<1$. In the Lorentz force ${\bi F}=e({\bi E}+{\bi u}\times{\bi B})$, if the fields are expressed in the international system of units, we need two conversion factors:
$K_{E}=\frac{e\hbar}{m^{2}c^{3}}=7.55676\cdot 10^{-19}\;{\rm m/V},$
$K_{B}=\frac{e\hbar}{m^{2}c^{2}}=2.26546\cdot 10^{-10}\;{\rm T}^{-1},$
such that the Lorentz force in natural units is
${\bi F}=K_{E}{\bi E}+K_{B}{\bi u}\times{\bi B},$
with $\bi E$} in V/m and $\bi B$} in teslas.
Therefore this system (19) and (20) of ordinary non-linear differential equations, written in natural units, depend on two physical dimensionless coefficients, $K_{E}E$ and $K_{B}B$, where $E$ and $B$ are, the intensity of the electric and magnetic field, respectively, in the I.S. of units, and all remaining variables are dimensionless. The stability of this non-linear system of ordinary differential equations depends on the values of these two coefficients and therefore sets a limit on the allowed values of the external fields.
The relationship between the International System of units and the Natural System of units has the following conversion factors:
IS to Natural System
Unit International System Natural System
$[L]$ length 1 m $2.58965\cdot 10^{12}\;$n.u.
$[M]$ mass 1 kg $1.09775\cdot 10^{30}\;$n.u.
$[T]$ time 1 s $7.76357\cdot 10^{20}\;$n.u.
$[Q]$ electric charge 1 C $6.24146\cdot 10^{18}\;$n.u.
$[I]$ current 1 A 0.008040710.00804071\;n.u.
$[E]$ electric field 1 V/m $7.55676\cdot 10^{-19}\;$n.u.
$[B]$ magnetic field 1 T $2.26546\cdot 10^{-10}\;$n.u.
[…]
4 Upper bound of the external fields
The acceleration of the CM (19) depends on the value of the external fields. The only constraint is that $v<1$. Let as assume that the electron is at rest under an external electric field. In a time $\Delta t$, $\Delta v=F\Delta t=K_{E}E\Delta t$. The electric field is bounded by
$E<\frac{1}{K_{E}\Delta t}=\frac{E_{S}}{\Delta t}.$ (21)
If the field is acting during one unit of natural time this implies that $E<E_{S}$ and thus the electric field is limited above. Similarly, under an external magnetic field $B<B_{S}$. The above upper limits of the external fields during a unit of natural time are called the Schwinger limits
$E_{S}=\frac{1}{K_{E}}$ $=$ $\frac{m^{2}c^{3}}{e\hbar}=1.323\cdot 10^{18}\;{\rm V/m},$ (22)
$B_{S}=\frac{1}{K_{B}}$ $=$ $\frac{m^{2}c^{2}}{e\hbar}=4.414\cdot 10^{9}\;{\rm T}.$ (23)
This limitation depends on the time $\Delta t$. The unit of natural time is $\tau_{0}=1.288\cdot 10^{-21}$s, and the time for a turn of the CC around the CM in the center of mass frame is $\pi$ in natural units. For a moving electron with a CM velocity $v$ this time for a complete cycle is $\gamma(v)\pi$.
If the transfer of energy from an electromagnetic field to matter is in the form of photons, as suggested by Planck’s hypothesis, we have to compute not only the modification of the CM velocity (which represents de transfer of energy and linear momentum) but also the modification of the angular momentum, such that the electromagnetic angular momentum transfer equals to $\hbar$, the spin of the photon. This will fix the necessary time $\Delta t$ to transfer a quantum of electromagnetic field. If the discrete time $\Delta t$ to transfer a quantum of spin $\hbar$ is computed, the allowed physical electromagnetic fields are bounded above by (21).
5 Mass and Spin invariance: Radiation reaction
According to the Atomic Principle, the mass $m$ and the absolute value of the spin in the center of mass frame $S(0)$ are the invariant properties of the electron that cannot be modified by any interaction. Therefore $dm/dt=0$, and mass is conserved. Taking the time derivative of $p^{\mu}p_{\mu}=H^{2}-{\bi p}^{2}=m^{2}$, we get
$H\frac{dH}{dt}-{\bi p}\cdot\frac{d{\bi p}}{dt}=0,$
and if we use (5) for the expressions of $H$ and $\bi p$} we arrive to
$\frac{dH}{dt}={\bi v}\cdot{\bi F},\quad dH=d{\bi q}\cdot{\bi F}.$
The variation of the mechanical energy of the electron is the work of the external Lorentz force along the CM trajectory. But the work of the external field acting on a point charge is the work along the CC trajectory, $d{\bi r}\cdot{\bi F}$. If these two trajectories are different, not all electromagnetic energy expended by the field is transformed into mechanical energy of the particle.
By energy conservation if this difference of energy ${\bi F}\cdot(d{\bi r}-d{\bi q})$, is positive the electron increases its mechanical energy, and the remaining part of this energy that has not been used in this acceleration process, is transformed into energy of the field. If this difference is negative, the electron loses mechanical energy and the difference is returned and transformed into energy of the field. It can be interpreted as if the particle radiates electromagnetic energy, because the total electromagnetic energy of the field has changed. If these two works are the same there is no radiation because all electromagnetic energy of this work is transformed into mechanical energy of the particle.
The possible radiation is related to the existence of an acceleration of the CM. The existence of an external force implies an acceleration of the CM, but it can happen that there exist CM accelerated motions without radiation.
The other invariant is that the absolute value of the spin in the center of mass frame $S(0)$ is conserved. In the preprint [15] it is analyzed this requirement of the conservation of the spin in the center of mass frame and this implies a modification of the dynamical equation (19) such that the breaking term ${\bi v}({\bi v}\cdot{\bi F})$ is replaced by the term ${\bi v}({\bi u}\cdot{\bi F})$, so that this differential equation becomes
$\frac{d{\bi v}}{dt}=\frac{1}{\gamma(v)}\left[{\bi F}-{\bi v}\left({\bi F}\cdot{\bi u}\right)\right].$ (24)
This corresponds to modify the dynamical equation $d{\bi p}/dt={\bi F}$, obtained from the Lagrangian formulation, by the equation
$\frac{d{\bi p}}{dt}={\bi F}-\gamma(v)^{2}\left[{\bi F}\cdot({\bi u}-{\bi v})\right]{\bi v}.$ (25)
The variation of the linear momentum contains a breaking term, opposite to the CM velocity $\bi v$}, that depends on the factor $\gamma(v)^{2}$ and the difference of the work per unit time of the external Lorentz force along the CC and the CM trajectories. This term can be interpreted as the reaction force produced by the interchange of photons between the particle and the field. In a continuous interaction, the term $\gamma(v)^{2}\left[{\bi F}\cdot({\bi u}-{\bi v})\right]d{\bi q}$, will be the linear momentum of the continuously radiated electromagnetic field in a time $dt$.
A single electron in a finite time $T$ radiates finite electromagnetic energy. This corresponds to the emission of a finite number of photons, let us say $N$. This means that when a single electron is accelerated it is not emitting continuous radiation but rather it emits on average a single photon after a discrete time $\Delta t=T/N$. Radiation is not a continuous process, so that we need to wait to a discrete time $\Delta t$ such that the quantum of energy corresponds to the transfer of angular momentum $\hbar$. The determination of this discrete time in different examples is left to a future research.
6 Conclusions
In the present formalism the electron is described by the evolution of a single point $\bi r$}, the center of charge. It is moving at the speed of light and satisfies a non-linear system of ordinary differential equations of order four. The center of mass $\bi q$}, can be defined simultaneously in every inertial frame in terms of $\bi r$} and some time derivatives. With this definition the dynamical equations can be decoupled into a system of second order ordinary differential equations for both points $\bi r$} and $\bi q$}. This implies that we have two different spins, the angular momentum with respect to $\bi q$}, ${\bi S}_{CM}$, and the angular momentum with respect to $\bi r$}, $\bi S$}. The spin ${\bi S}_{CM}$ is conserved in the free motion but the spin $\bi S$} satisfies the same dynamical equation than Dirac’s spin operator in the quantum case.
The existence of two different points, one the CC $\bi r$}, associated to the electromagnetic structure, and another the CM $\bi q$} with the inertia of the particle, implies that when it interacts with an external electromagnetic field, there is a difference between the energy expended by the field and the mechanical energy acquired by the particle. This difference implies that the particle radiates. This modifies the dynamical equations in order that the two intrinsic properties of the particle, the mass $m$ and the absolute value of the spin $S(0)$ remain unchanged under an external interaction. The dynamical equations are modified by a radiation-reaction force, opposite to the CM velocity.
Radiation is not a continuous process. Nevertheless the classical description of the electron and Maxwell’s equations are based on continuous and differentiable variables and fields. If according to Planck’s hypothesis radiation is the emission of photons, these are not emitted continuously, and a discrete time must exist between two consecutive emission of photons by any single electron, until the continuously radiated electromagnetic angular momentum amounts the quantum $\hbar$.
The complete set of dynamical equations can be expressed in terms of dimensionless variables or natural units, where the only physical parameters that characterize the system of equations are the intensity of the electric and magnetic fields in natural units, so that the stability of these equations will be restricted by these values. The allowed electric and magnetic fields are not unbounded. The existence of this discrete time $\Delta t$ between the emission of consecutive photons implies that the allowed values of the electric and magnetic field are bounded above if the dynamical equations have to remain stable. For greater values of the fields the constraint that the CM velocity $v<1$, does not hold.